OpenStudy (twalt13):

please can anyone help me ? Write the sum using summation notation, assuming the suggested pattern continues. -10 - 2 + 6 + 14 + ... + 110

OpenStudy (anonymous):

\[d=8, a _{n}=110 \] \[a _{n}=a _{1} +d(n-1)\] \[a _{1}=-10 \] 110=-10+8(n-1) 120=8(n-1) n-1=15 n=16

OpenStudy (anonymous):

110 is 16th term

OpenStudy (twalt13):

ima show you my answer choices

OpenStudy (anonymous):

\[_{Sn}=\frac{ n }{ 2 }(2a _{1}+(n-1)d)\] then substitute the above values \[S _{16}=8(-10+110)=8*100=800\]

OpenStudy (anonymous):

the answer is 800

OpenStudy (twalt13):

\[\sum_{n=0}^{15}(-10+8n) \sum_{n=0}^{\infty}(-10+8n) \sum_{n=0}^{15}-80n \sum_{n=0}^{\infty}-80n\]

OpenStudy (twalt13):

@getusel which one would be the choice ?

OpenStudy (anonymous):

the first one is right

OpenStudy (twalt13):

thanks

OpenStudy (anonymous):

\[\sum_{0}^{15}(-10+8n)\]

OpenStudy (anonymous):

it is my pleasure

OpenStudy (twalt13):

@getusel i have another problem im trying to solve